rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#11058 — FS#14973 — Incoming calls

Attached to Project— VoIP
Incident
Backend / Core
CLOSED
100%
An incident regarding the History is in progress on 3 of our interconnections.
Incoming calls on OVH network have been impacted.
We will keep you informes on the development of this taso.
Date:  Tuesday, 06 March 2018, 09:31AM
Reason for closing:  Done
Comment by OVH - Tuesday, 13 October 2015, 16:25PM

We are in touch with the History operator in order to solve the incident.
The input traffic has been secunde by 2 of our other interconnections.


Comment by OVH - Tuesday, 13 October 2015, 16:27PM

Expert work is in progress with the History operator.

3 of our 4 interconnections havebeen impacted.


Comment by OVH - Tuesday, 13 October 2015, 16:28PM

The problem was identified by the operator's teams.

It affects the area of our datacentres.

We are making every effort to solve it as soon as possible.


Comment by OVH - Tuesday, 13 October 2015, 16:29PM

The operator has just started to work on site.


Comment by OVH - Tuesday, 13 October 2015, 16:44PM

The theft of copper pairs in the neighborhood paris 19
is causing the problem. We have 4 interconnections between
Orange and OVH, 3 on copper and 1 on on the
optical fiber. The optical fiber interconnection is UP and those of copper were vandalized.
There is a lack of capacity 30% to take all incoming calls. By 5:30 pm, the number of
calls will be low enough to be cashed in degraded mode.

Orange is operating on site in order to to replace the copper pairs necessary for interconnection. It will take
a few hours.

In parallel, a dual IP interconnection between Orange and OVH is being created, one interconnection towards TH2 and
the other towards GSW.


Comment by OVH - Tuesday, 13 October 2015, 16:45PM

The call peak falls, the number of calls is low enough to be cashed.

The operator is still working on site.


Comment by OVH - Tuesday, 13 October 2015, 16:46PM

Orange has yet to work tomorrow morning to return the situation back to normal.
The current capacity is more than enough to keep the load of the weekend.


Comment by OVH - Tuesday, 13 October 2015, 16:47PM

We are expecting the operator's feedback on the work progress.


Comment by OVH - Tuesday, 13 October 2015, 17:00PM

We have 4 links to the 3 different Orange infrastructures. 3 PRO in Paris. every POP is independent and provides routing calls to our infrastructure.
When a link is DOWN, the Orange PRO should return the call to another PRO and so on until the call ends on the PRO that is connected to our infrastructure.

We have noticed that this is not working well and some calls are not successful. It is possible that the call arrives on one of the three PRO when it is
DOWN, so it returns the call to another PRO. If this 2nd PRO is also DOWN and then returns the call to the next PRO.
Except it must return the call to the first that PRO, which will cut the call off becuse it has already received it.
In short, our configuration works fine with one PRO (of 3) down, but not with 2 PROs down.


Comment by OVH - Tuesday, 13 October 2015, 17:03PM

The operation is still in progress following the act of vandalism.
Orange indicates that the problem is very important and that these teams are working on it.
ETA is not specified yet. Regular monitoring is scheduled for this afternoon.


Comment by OVH - Tuesday, 13 October 2015, 17:11PM

The outcome of the intervention is not good.
The rewiring work is much larger than expected and the estimate of the resolution is postponed to the middle of next week (!).
The next review will be stated on Monday.

We are studying several alternatives in cooperation with the History operator to handle the load of incoming calls of the week.
Among others, interconnection switching in order to listen to traffic to OVH through SFR with which we have
a lot of capacity.
We will inform you of the actions taken upon confirmation on Monday morning.